undeadknight26 (undeadknight26):

Donna teaches dance in groups of five students. She rents a hall for the dance classes at $450 per month. She charges $40 per hour per student. Donna teaches four groups of students, and her net earnings per month are $4,350. Part A: If each student learns dance for the same number of hours, write an equation that can be used to calculate the number of hours each student learns dance in a month. (5 points)

undeadknight26 (undeadknight26):

Just need an equation i should be able to solve it.

undeadknight26 (undeadknight26):

@SolomonZelman

OpenStudy (solomonzelman):

Can you write out the equations for this question?

OpenStudy (solomonzelman):

\[student~\times~hours~\times~40=4350+350=4800\] +350 because even when she pays 350 she gets 4350.

OpenStudy (solomonzelman):

\[hours~~=~~4800~~\div~~40~~=~~120\]

OpenStudy (solomonzelman):

\[120~\div~5~=~24\]because there are 5 students.

undeadknight26 (undeadknight26):

oh ok thx! but why does she have to pay 350?

OpenStudy (solomonzelman):

oh it's +450, not 350.

OpenStudy (solomonzelman):

but the answer is same.

undeadknight26 (undeadknight26):

oh ok.

undeadknight26 (undeadknight26):

and arent there 20 students because 5 x 4?

OpenStudy (solomonzelman):

\[4350+450=4800\]\[4800~\div~40=~120\]\[120~/~5~=~24\]

OpenStudy (solomonzelman):

4 groups?

OpenStudy (solomonzelman):

All I know it is 24 hours per group.

OpenStudy (solomonzelman):

in a month.

OpenStudy (solomonzelman):

To find the number of hours per student, you would need to divide 24 by number of students. the PROBLEM is that the number of students is NOT GIVEN.

undeadknight26 (undeadknight26):

Yes there are 4 groups so that means the answer is each student works for 6 hours?

OpenStudy (solomonzelman):

You aren't given the number of students per group though, so I can only say 24h per group each month.

undeadknight26 (undeadknight26):

Donna teaches dance in groups of five students. and later they say there are 4 groups :p heheh

undeadknight26 (undeadknight26):

can you help me with a few more?

OpenStudy (solomonzelman):

I might be able to, but my sister has an engagement party and I have to always leave and help at different times. (Not incl. other circumstances)

undeadknight26 (undeadknight26):

lol well tell her that the undead wishes her a good luck...it'll scare the crab out of here if you say it creepily...

OpenStudy (solomonzelman):

I don't want to scare her like that....

OpenStudy (solomonzelman):

She would be like "wtf" (in a different form of course) but ty for a good wish.

undeadknight26 (undeadknight26):

hehe your too good of a person...

OpenStudy (solomonzelman):

Me? nice to hear that some people think I am nice (as opposed to what my little brother thinks when he steals my $300 headphones a 10th time...)

undeadknight26 (undeadknight26):

lol ahahah...im a pranker and i say a opportunity like that only comes once in a life time...lol lucky 300 dollar headphones...i got 15 dollar ones hahah..